OpenStudy (anonymous):

Solve algebraically for x: 3(x+1) - 5x = 12 - (6x-7)

OpenStudy (johnweldon1993):

so distribute the 3 into those parenthesis...and also distribute the '-' into those parenthesis

OpenStudy (anonymous):

Okay

OpenStudy (johnweldon1993):

and what do you get from that?

OpenStudy (anonymous):

3x +1

OpenStudy (anonymous):

I mean 3x + 3

OpenStudy (johnweldon1993):

There you go.....right... 3x + 3 - 5x = 12 - (6x - 7) we can simplify the left hand side for now 3x - 5x is?

OpenStudy (anonymous):

-2x

OpenStudy (johnweldon1993):

alright...so altogether we have -2x + 3 = 12 - (6x - 7) now on to the right hand side after distributing the '-' into the parenthesis what do you have?

OpenStudy (anonymous):

Ohhhh I see what I did I didn't distribute the right hand side thanks!

OpenStudy (johnweldon1993):

no problem!

OpenStudy (johnweldon1993):

what do you get for your final 'x'...?

OpenStudy (anonymous):

-4?

OpenStudy (johnweldon1993):

'-' 4 ? hmm...lets do it out -2x + 3 = 12 - 6x + 7 -2x + 3 = 19 - 6x +6x +6x 4x + 3 = 19 -3 -3 4x = 16 --------- 4 4 x = 4